The 17th term of an A.P is 5 more than twice it's 8th term. If the 11th term of the A.P Is 43 find the math term. 

Grade:10

2 Answers

Arun
4 years ago
a₁₇ = 5 + 2a₈                  and a₁₁ = 43 
we know nth of an A.P.  =  a + (n-1) d    where  a = first term 
                                                                           d = common difference of A.P
                                                  
         therefore
           a + (17  - 1) d  =  5 + 2{ a + (8 - 1 ) d}  
           a + 16 d =  5 + 2a + 14d        2d - a = 5        .......(1)

also      a + (11 - 1)d = 43              a + 10d = 43        .........(2)

solving 1 and 2 gives        12d = 48              d = 4 
                     a =  3
hence nth tern would be       =  3 + (n-1)4  ans 

 
Vikas TU
4 years ago
Dear student 
Please specify in the question what is math term ? You are asking abouth m th term or math term 
Regard 
Vikas